expect to retire You are 35 years old today and are considering your retirement needs. You at age 65 and your actuarial tables suggest that you will live to be 100. You want to move to the Bahamas when you retire. You estimate that it will cost you $ 300,000 to make the move (on your 65th birthday) and that your living expenses will be $30,000 a year (starting at the end of year 66 and continuing through the end of year 100) after that. a. How much will you need to have saved by your retirement date to be able to afford this 18. course of action? b. You already have $50,000 in savings. If you can invest money, tax-free, at 8% a year, how much would you need to save each year for the next 30 years to be able to afford this retirement plan? c. If you did not have any current savings and do not expect to be able to start saving money for the next 5 years, how much would you have to set aside each year after that to be able to afford this retirement plan?
Answer #1

Answer 18 a:

Funds required as on Retirement date (65th Birthday):

(i) Fund required on 65th birthday to make a move Bahamas = $300,000

(ii) Living expense of $30,000 a year starting end of year 66 to end of year 100

Interest rate = 8%

To calculate PV as on retirement date we will use PV function of excel

PV (rate, nper, pmt, fv, type)

= PV (8%, 35, -30000, 0, 0)

= $349,637.0465

Funds required as on Retirement date = $300,000 + $349,637.05 = $649,637.05

Funds required as on Retirement date = $649,637.05

Answer 18 b:

At age 35:

Savings = $50,000

Funds required on retirement date (65th birthday) = $649,637.05

To get required yearly saving we will use PMT function of excel:

PMT (rate, nper, pv, fv, type)

PMT (8%, 30, 50000, -649637.05, 0)

=$1,293.26

Annual saving required = $1,293.26

Answer c:

FV = PV (1 + Interest rate) Number of years

Future value of 50,000 savings = $50,000 * (1 + 8%) 30 = $503,132.844 =$503,132.84

Balance fund required as of retirement date from annual savings = $649,637.0465 - $503,132.844 = $1,46,504.20

Annual savings required after 5 years:

= PMT (8%, 25, 0, -146504.20, 0)

= $2004

Annual savings required after 5 years = $2004